OpenStudy (anonymous):

A ball is thrown vertically upward.After (t) seconds, its height (h) (in feet) is given by this function h(t)=96t-16t^2. What is the maximum height that the ball will reach? Check Answer for me please?

OpenStudy (anonymous):

I got the answer 144 ft. Is it correct?

OpenStudy (freckles):

Calculus or algebra ? (we can do either here)

OpenStudy (anonymous):

College Algebra

OpenStudy (freckles):

Ok so I assumed you found the x-coordinate of the vertex and then plugged that into the function to find the max height?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

can I ask what you got for the x-coordinate of the vertex?

OpenStudy (anonymous):

3

OpenStudy (freckles):

awesome

OpenStudy (freckles):

let me check the rest one sec

OpenStudy (freckles):

\[h(3)=96(3)-16(3)^2=288-144=144\] you are right

OpenStudy (anonymous):

okay thankyou :)

OpenStudy (freckles):

you get a medal because you answered your own question

OpenStudy (freckles):

good job :)

OpenStudy (anonymous):

awesome :)